OpenStudy (anonymous):

power series solution y' + y/x = 0

myininaya (myininaya):

xy'+y=0 (xy)'=0 xy=C y=C/x

OpenStudy (anonymous):

\[dy/dx + y/x = 0\] \[dy/dx = -y/x\] \[dy/y = -dx/x\] \[\int\limits(dy/y) = \int\limits(-dx/x)\] \[lny = -lnx + lnk\] \[lny = -\ln(x/k)\] \[e ^{lny} = e^{-\ln(x/k)}\] \[y = k/x\], where k is an arbitrary constant.